### v1

@@ 4.1 Quantum convolutional autoencoder (QCAE)
In this study, we introduce the QCAE architecture designed for image denoising. Leveraging the CCAE model as its foundation, the QCAE incorporates a QAOA circuit to replace the CCAE’s latent space. The QCAE process involves several key stages: i) initializing the input image data, ii) introducing noise to the input image data to generate a noisy input, iii) encoding the input image data into a quantum state, iv) using a QAOA circuit to manipulate the quantum state, v) measuring the output of the QAOA circuit, and vi) generating denoised input image data through a classical decoder based on the QAOA circuit measurement results [17]. Fig. 3 showcases the QCAE architecture and the steps involved in image denoising. Given a random noise-generating function $S$, the process involves taking an input data $x$ and subjecting it to the influence of $S$ to produce a noisy input image data version represented as $S(x)$. Subsequently, this noisy input $S(x)$ serves as the input for the QCAE. A proficient QCAE model should be capable of efficiently handling noisy images, adeptly processing them to reconstruct the original image, all the while eliminating any undesirable noise.
@@
In the QCAE, the convolutional encoder is responsible for extracting and compressing crucial features from the input image data. Through a carefully designed sequence of convolutional layers, pooling layers, and activation functions, the encoder systematically reduces the dimensionality of the input image data. This process results in the creation of a condensed, low-dimensional latent representation that encapsulates the essential features of the input image [7, 41]. By replacing the classical latent space with the QAOA circuit, one can obtain latent representations that are more effectively optimized for preserving the most crucial features within the input image data [35, 22]. In contrast to the convolutional encoder, the convolutional decoder consists of a series of transposed convolutional layers, upsampling layers, and activation functions. Transposed convolutional layers essentially perform the inverse operation of convolutional layers. These layers augment the spatial dimensions of the feature maps, simultaneously acquiring the ability to reconstruct lost details of the encoded input image data.
@@ 4.2 Quantum approximate optimization algorithm (QAOA)
QAOA, originally proposed by Farhi et al. [13], is a heuristic approach specifically designed to tackle the maxcut problem. It belongs to the group of hybrid algorithms and requires, in addition to executing shallow quantum circuits, a classical optimization process to optimize the parameters and improve the quantum circuit itself [16, 8]. The performance of QAOA continually improves with increasing values of $p$, contributing to the achievement of so-called quantum supremacy [14]. As depicted in Fig. 4, the maximum or minimum value of objective function can be obtained by iteratively finding the optimal values for the parameters $\gamma$ and $\beta$. The iterative process for determining the optimal parameters relies on classical optimization methods. In the context of QCAE, the latent space aims to acquire a compressed representation of the input image data, capturing inherent patterns and relationships among the data points. This representation aims to be more compact than the original input data while retaining essential information. Given that the primary objective of QAOA is to identify the optimal solution for a given problem, the QAOA circuit is utilized to discern the quantum information of the input state through the optimal latent representation.
@@
Here, we then introduce the process of injecting parameter values into rotation gates $R$ parameter placeholders. The training process of the QAOA circuit begins with input image data $x$ and a set of rotation gates $R$ at level $p$. First, the input image data are encoded using a classical encoder $\vec{x}=E(x)$, where $\vec{x}$ is an encoded input image data. Then, this data undergoes additional classical processing through convolutional layers, pooling layers, and is transformed into a linear form of $y$ to the size of $2p$, where $p\geq 1$ is the QAOA circuit level. Classical normalization and optimization techniques are applied to the variable $y$ to fine-tune the parameters, aiming to minimize the objective function (cost function). Fig. 5 illustrates the schematic illustrates the classical encoder in encoding and optimizing input image data for the QAOA circuit at any given $p$ levels.
@@ 4.3 Parameter-shift rule optimization
In this section, we present a PSR optimization method inspired by [47]. The PSR is a promising approach for evaluating the gradients of QAOA circuits on real quantum machines [10, 49]. Gradients derived from these quantum circuits significantly contribute to the optimization phase of VQAs. In the context of hybrid quantum-classical algorithms, we first construct a quantum circuit and subsequently adjust its parameters to minimize an objective function or cost function [36]. In our proposed method, we compute the gradient of each parameter in the QAOA circuit by per the parameter and determining the difference between two forming a double shift, positive shift and negative shift, on outputs, without modifying the QAOA circuit structure or using any ancilla qubits [47]. Suppose we have an $n$-qubit QAOA circuit which is parameterized by $m$ parameters, $\theta=[\theta_{1},\dotsm,\theta_{i},\dotsm,\theta_{m}]$, the objective function $f(\theta)$ of a quantum circuit can be represented by a circuit function as follows:
@@
where $H$ is the Hermitian generator of $U(\theta_{i})$ with only two unique eigenvalues $+1$ and $-1$. In this way, the differences in the gradients of the circuit function $f$ with respect to $\theta_{i}$ are:
$\begin{array}[]{l}\frac{\partial f(\theta)}{\partial\theta_{i}}=\frac{1}{2}\Bigl{(}f(\theta_{+})-f(\theta_{-})\Bigl{)},\\ \\ \theta_{+}=[\theta_{1},\dotsm,\theta_{i}+\frac{\pi}{2},\dotsm,\theta_{m}],\\ \\ \theta_{-}=[\theta_{1},\dotsm,\theta_{i}-\frac{\pi}{2},\dotsm,\theta_{m}],\end{array}$ (20)
@@
Fig. 6 visually depicts the procedural steps of the technique PSR. In each iteration, we perform dual shifts of the parameter $\theta_{i}$ by $+\frac{\pi}{2}$ and $-\frac{\pi}{2}$, respectively. After each parameter shift, we execute the shifted QAOA circuit on either a local quantum simulator or a quantum device. Upon obtaining the results from the two shifted circuits, we apply Equation 20 to calculate the upstream gradient. Following the PSR, we then execute the QAOA circuit without any parameter shift and record the measurement results. Subsequently, we apply softmax and cross-entropy functions to the obtained logits, ultimately producing training loss. The backpropagation process is initiated solely from the loss to the logits, generating downstream gradients. Finally, the gradient is computed by taking the dot product between the upstream and downstream gradients, as outlined in Eq. 23, resulting in the final gradient.
@@ 5.1 Datasets
We perform both CCAE and QCAE on the well-known MNIST dataset. This dataset comprises handwritten digits presented as grayscale images. The MNIST dataset comprises 60,000 samples in the training set and 10,000 samples in the testing set, each represented by pixels arranged in a shape of $28\times 28\times 1$. To incorporate noise into the datasets, we used a function that adds random Gaussian noise to the images. The original and noisy images are visualized in Fig. 3. For the MNIST dataset, we selected two specific classes from MNIST (0 and 1). For a fair comparison between CCAE and QCAE, the training datasets comprise the initial 2000 samples.
@@ 5.2 Environments
The experiments were conducted using the PyTorch [31] and Qiskit [34] frameworks. The proposed QCAE is tested on both a local quantum simulator equipped with a quantum assembly (QASM) simulation and a 27-qubit IBMQ Mumbai quantum machine. For noisy simulators, we leverage the noise model from IBM Mumbai and inject it into the QASM simulator. For notation simplicity, we designate the QASM simulator as the noiseless simulator, the noisy QASM simulator as the noisy simulator, and the actual quantum machine as the real machine.
@@
Figure 8: (Left) Comparison of training loss between the CCAE and QCAE models showcases a decrease in loss as the number of epochs increases. (Right) Comparison of SSIM between CCAE and QCAE shows an average SSIM of approximately 0.45 and 0.75, respectively.
@@ 6.1 Training of the CCAE and QCAE
In this evaluation, two binary classes were extracted from the MNIST dataset (0 and 1). To gain insights into the distinctions between denoised samples generated by both the CCAE and QCAE models and the original samples, we employ a metric called the structural similarity index measure (SSIM) for analyzing image quality [44], ranging from 0 (worst) to 1 (best). To illustrate the training performance, the left side of Fig. 8 displays the training loss. Our proposed QCAE consistently outperforms CCAE, achieving a lower training loss. The right side of Fig. 8 illustrates the SSIM values over 50 training epochs, demonstrating that QCAE generates denoised images closely resembling and showcasing higher quality compared to those produced by CCAE. This distinction is particularly evident when considering the average of SSIM 0.45 for CCAE and SSIM 0.75 for QCAE. These findings highlight that QCAE consistently delivers robust results when employed for denoising images compared with CCAE.
Fig. 9 depicts original images alongside those with added noise, where high corruption levels render them barely visible to the human eye. The denoising results generated using both QCAE and CCAE models are shown alongside the noisy and original images. Notably, in the third row and middle column, the CCAE struggles to fully denoise or reconstruct the image, highlighting a limitation. Furthermore, CCAE faces challenges in effectively denoising, resulting in denoised results that exhibit a subtle blurring effect when compared to the sharper outcomes achieved by QCAE, with an average SSIM for CCAE at 0.45 and the SSIM for QCAE at 0.75. This performance gap emphasizes the effectiveness of our proposed QCAE, which outperforms CCAE as its classical counterpart.
Figure 9: Denoising results on MNIST datasets. The top row displays original images, and the second row shows their noisy counterparts. The third row showcases images denoised with CCAE, and the last row features images denoised using the proposed QCAE method.
@@ 6.2 Denoising performance
Here, we conducted denoising of the MNIST images across three distinct scenarios: noiseless, noisy, and real. To effectively demonstrate the performance of the QCAE, we introduced variations in Gaussian noise factors applied to the images because the MNIST does not carry noise. For each scenario, we experiment with the Gaussian noise factor $\sigma=\{0.25,0.5,0.75,1\}$, and choose the optimal combination from 1 out of 12 options based on the reasonable noise factor and highest SSIM for further experiments. The QCAE performance is evaluated for the denoising capability in both quantum simulations and actual machines.
Fig. 10 illustrates the performance of CCAE and specifically QCAE when tested on noisy images with varying noise factors. The findings reveal a notable reduction in SSIM value as the noise factor increases in the image. However, even in the challenging conditions of current noisy and real environments, our proposed QCAE showcases results comparable to noiseless environments. This underscores the robustness of our QCAE, confirming its ability to operate effectively in diverse quantum computer environments. One should take note that the noisy configuration outperforms the real setting, even though both utilize the same IBM machine. This superiority arises from the fact that the noise model incorporates various critical factors, including basic gate error, gate length, decoherence times, and the readout error probability of each qubit. These parameters are calibrated on a daily basis, contributing to a slightly lower level of noise precision compared to the real configuration [24].
Figure 10: Comparative analysis of the CCAE and specifically QCAE performance under varying conditions, including noiseless, noisy, and real scenarios, with noise factor variations represented by $\sigma=\{0.25,0.5,0.75,1\}$.
To analyze the extent of QCAE denoising capabilities, we utilize it to restructure, denoise, and generate images from a set of noisy inputs with varying noise factors values. Even in the presence of high levels of noise, characterized by a noise factor of $\sigma=1$, images that appear indistinguishable to the human eye are effectively processed by QCAE. The algorithm consistently delivers accurate and visually appealing results, demonstrating its robust performance in handling noisy images. Despite testing in noisy and real quantum machine environments, the image maintains readability and closely resembles the original, although with a slight degradation compared to a noiseless environment. Denoising results, along with variations in Gaussian noise levels and quantum machine environments, are depicted in Fig. 11.
@@ 6.3 Comparison of PQCs and PSR optimization
Figure 12: The SSIM value of $p$ ranges from 1 to 5 across all circuits, indicating that better solutions are associated with longer circuit layers.
We also conducted comprehensive experiments on all circuits, both with and without PSR, resulting in a noteworthy reduction in SSIM values when tested without PSR, as shown in Table 1. Remarkably, our proposed circuit demonstrated the ability to draw comparisons with other circuits, even in the absence of PSR. One plausible explanation is that $R_{Z}$ gates can be implemented with greater efficiency compared to other single-qubit rotation operations [25]. As we observed, the circuits exhibit lower and significantly higher SSIM values with and without PSR, respectively. This underscores the significance of incorporating PSR optimization to enhance performance. It is noteworthy that the $p$ value significantly influences the results, underscoring its high impact on the outcomes. However, as the parameter $p$ increases, there is a corresponding growth in the circuit execution time. Additionally, our proposed method requires a minimum of two qubits for the efficient denoising of images. It also illustrates the progressive increase in SSIM values corresponding to the varying number of $p$ layers. As additional layers are added, the QAOA showcases the potential for enhanced approximations towards the optimal solution. Furthermore, it presents the results of experiments conducted on QAOA, both with and without PSR optimization. The findings emphasize the significance of employing this optimization technique to identify a lower-cost function, leading to higher accuracy. Through the utilization of the PSR optimization, there is a notable enhancement in SSIM value. This enhancement showcases an improvement of up to 35% when compared with non-optimized conditions.

### v2

@@ 4.1 Quantum convolutional autoencoder (QCAE)
In this study, we introduce the QCAE architecture designed for image denoising. Leveraging the CCAE model as its foundation, the QCAE incorporates a QAOA circuit to replace the CCAE’s latent space. The QCAE process involves several key stages: i) pre-processing the input image data, which includes normalizing pixel values to the [0, 1] range, ii) introducing Gaussian noise to the input image data to generate a noisy input, iii) transforming the noisy input through classical convolutional layers to produce a latent representation, which is then encoded into a quantum state, iv) using a QAOA circuit to manipulate the quantum-encoded latent vector, v) measuring the output of the QAOA circuit, and vi) generating denoised input image data through a classical decoder based on the QAOA circuit measurement results [29]. Fig. 3 showcases the QCAE architecture and the steps involved in image denoising. Given a random Gaussian noise-generating function $S$, the process involves taking an input data $x$ and subjecting it to the influence of $S$ to produce a noisy input image data version represented as $S(x)$. Subsequently, this noisy input $S(x)$ serves as the input for the QCAE.
@@
The training process of the QAOA circuit begins with input image data $x$ and a set of rotation gates at level $p$. First, the input image is encoded by a classical encoder $\vec{x}=E(x)$, where $\vec{x}$ represents the encoded latent vector [14, 30]. This latent representation is further compressed into a linear form $y$ with dimension $2^{p}$, where $p\geq 1$ corresponds to the QAOA circuit depth. While this step reduces the dimensionality classically, the QAOA circuit operates in a quantum state space of dimension $2^{n}$, with $n$ denoting the number of qubits, thereby enabling access to richer high-dimensional representations [15, 31]. Classical normalization and optimization methods are then applied to fine-tune the parameters and minimize the cost function. Fig. 5 illustrates this workflow, showing how the classical encoder prepares input data and optimizes it for QAOA circuits at various depths $p$.
@@ 4.2 Quantum approximate optimization algorithm (QAOA)
QAOA, originally proposed by Farhi et al. [32], is a heuristic approach specifically designed to tackle the maxcut problem. It belongs to the group of hybrid algorithms and requires, in addition to executing shallow quantum circuits, a classical optimization process to optimize the parameters and improve the quantum circuit itself [33, 34]. The performance of QAOA continually improves with increasing values of p, where p denotes the number of alternating layers (depth) in the QAOA circuit, which controls the expressiveness of the variational ansatz [35]. As depicted in Fig. 4, the maximum or minimum value of the objective function can be obtained by iteratively finding the optimal values for the parameters $\gamma$ and $\beta$. The iterative process for determining the optimal parameters relies on classical optimization methods. In the context of QCAE, QAOA is adapted to create a quantum latent space representation. This latent space aims to generate a compressed encoding of input image data, capturing essential patterns and relationships while reducing dimensionality. The QAOA circuit in QCAE leverages its optimization capabilities to discern quantum information from the input state, creating an optimal latent representation. This approach utilizes QAOA’s problem-solving framework to efficiently encode input data in a quantum latent space, potentially offering advantages in data compression and feature extraction for image processing tasks.
@@
Here, we then introduce the process of injecting parameter values into rotation gates $R$ parameter placeholders. The training process of the QAOA circuit begins with input image data $x$ and a set of rotation gates $R$ at level $p$. First, the input image data are encoded using a classical encoder $\vec{x}=E(x)$, where $\vec{x}$ is an encoded input image data. Then, this data undergoes additional classical processing through convolutional layers, pooling layers, and is transformed into a linear form of $y$ to the size of $2p$, where $p\geq 1$ is the QAOA circuit level. While $y$ is classically compressed, the quantum circuit operates on a state space of dimension $2^{n}$, where $n$ is the number of qubits, providing access to high-dimensional quantum representations Classical normalization and optimization techniques are applied to the variable $y$ to fine-tune the parameters, aiming to minimize the objective function (cost function). Fig. 5 illustrates the schematic illustrates the classical encoder in encoding and optimizing input image data for the QAOA circuit at any given $p$ levels.
@@ 4.3 Parameter-shift rule optimization
In this section, we present a PSR optimization method inspired by [36]. The PSR is a promising approach for evaluating the gradients of QAOA circuits on real quantum machines [37, 38]. Gradients derived from these quantum circuits significantly contribute to the optimization phase of VQAs. In the context of hybrid quantum-classical algorithms, we first construct a quantum circuit and subsequently adjust its parameters to minimize an objective function or cost function [39]. In our proposed method, we compute the gradient of each parameter in the QAOA circuit by per the parameter and determining the difference between two forming a double shift, positive shift and negative shift, on outputs, without modifying the QAOA circuit structure or using any ancilla qubits [36]. Suppose we have an $n$-qubit QAOA circuit which is parameterized by $m$ parameters, $\theta=[\theta_{1},\ldots,\theta_{i},\ldots,\theta_{m}]$, the objective function $f(\theta)$ of a quantum circuit can be represented by a circuit function as follows:
@@
where $H$ is the Hermitian generator of $U(\theta_{i})$ with only two unique eigenvalues +1+1 and 1-1. In this way, the differences in the gradients of the circuit function $f$ with respect to $\theta_{i}$ are:
$\begin{array}[]{l}\displaystyle\frac{\partial f(\theta)}{\partial\theta_{i}}=\frac{1}{2}\!\left(f(\theta_{+})-f(\theta_{-})\right),\\[8.0pt] \theta_{+}=[\,\theta_{1},\,\ldots,\,\theta_{i}+\frac{\pi}{2},\,\ldots,\,\theta_{m}\,],\\[6.0pt] \theta_{-}=[\,\theta_{1},\,\ldots,\,\theta_{i}-\frac{\pi}{2},\,\ldots,\,\theta_{m}\,],\end{array}$ (20)
@@
Fig. 6 visually depicts the procedural steps of the technique PSR. In each iteration, we perform dual shifts of the parameter $\theta_{i}$ by $\frac{\pi}{2}$ and $\frac{\pi}{2}$, respectively. After each parameter shift, we execute the shifted QAOA circuit on either a local quantum simulator or a quantum device. Upon obtaining the results from the two shifted circuits, we apply Equation 20 to calculate the upstream gradient. Following the PSR, we then execute the QAOA circuit without any parameter shift and record the measurement results. Subsequently, we apply softmax and cross-entropy functions to the obtained logits, ultimately producing training loss. The backpropagation process is initiated solely from the loss to the logits, generating downstream gradients. Finally, the gradient is computed by taking the dot product between the upstream and downstream gradients, as outlined in Eq. 23, resulting in the final gradient.
@@ 5.1 Datasets
We perform both CCAE and QCAE on the well-known MNIST dataset. This dataset comprises handwritten digits presented as grayscale images. The MNIST dataset comprises 60,000 samples in the training set and 10,000 samples in the testing set, each represented by pixels arranged in a shape of $28\times 28\times 1$. To incorporate Gaussian noise into the datasets, we used a function that adds random Gaussian noise with standard deviation values $\sigma=\{0.25,0.5,0.75,1\}$ to the images. The original and noisy images are visualized in Fig. 3. Although restricting MNIST to digits 0 and 1 yields 12,000 training and 2,000 test samples, in this study we used smaller subsets to reduce computational cost and accommodate quantum hardware limitations. Specifically, we selected the first 2,000 samples of digits 0 and 1 from the training set (total 4,000 training images) and the first 200 samples of digits 0 and 1 from the test set (total 400 test images). Gaussian noise was applied only to the training set during model optimization, while the test set remained unseen throughout training and was used exclusively for evaluation. This ensures that no overlap occurs between training and testing data. A separate validation set was not used during training due to the dataset size; model performance was assessed directly on the test set after training. While MNIST is widely used for benchmarking, it has limitations in demonstrating the generalizability and real-world applicability of machine learning models, particularly in the quantum domain, as outlined in [43].
@@ 5.2 Environments
The experiments were conducted using the PyTorch [40] and Qiskit [44] frameworks. The proposed QCAE is tested on both a local quantum simulator equipped with a quantum assembly (QASM) simulation and a 27-qubit IBMQ Mumbai quantum machine. For noisy simulators, we leverage the quantum noise model from IBM Mumbai and inject it into the QASM simulator. For notation simplicity, we designate the QASM simulator as the noiseless simulator, the noisy QASM simulator as the noisy simulator, and the actual quantum machine as the real machine.
@@
Figure 8: Training loss comparison between CCAE and QCAE over 50 epochs.
@@ 6.1 Training of the CCAE and QCAE
In this evaluation, two binary classes were extracted from the MNIST dataset (0 and 1). To gain insights into the distinctions between denoised samples generated by both the CCAE and QCAE models and the original samples, we employ a metric called the structural similarity index measure (SSIM) for analyzing image quality [46], ranging from 0 (worst) to 1 (best). Fig. 8 illustrates the training loss over 50 epochs, where QCAE consistently outperforms CCAE. The corresponding SSIM values are presented in Fig. 10 across all scenarios (Noiseless, Noisy, and Real). QCAE achieves higher SSIM values than CCAE in most cases. However, at higher noise levels ($\sigma=0.75$ and $1$) in the Real and Noisy scenarios, its SSIM values are slightly lower than those of CCAE, reflecting the impact of quantum hardware noise. Overall, these findings highlight that QCAE consistently delivers robust results when employed for denoising images compared with CCAE.
Fig. 9 depicts original images alongside those with added Gaussian noise, where high corruption levels render them barely visible to the human eye. The denoising results generated using both QCAE and CCAE models are shown alongside the noisy and original images. Notably, in the third row and second column, the CCAE struggles to fully denoise or reconstruct the image, highlighting a limitation. Furthermore, CCAE faces challenges in effectively denoising, resulting in denoised results that exhibit a subtle blurring effect when compared to the sharper outcomes achieved by QCAE, with an average SSIM for CCAE at 0.45 and the SSIM for QCAE at 0.75. This performance gap emphasizes the effectiveness of our proposed QCAE, which outperforms CCAE as its classical counterpart.
Figure 9: Denoising results on MNIST datasets. The top row displays original images, and the second row shows their noisy counterparts. The third row showcases images denoised with CCAE, and the last row features images denoised using the proposed QCAE method. All images are from the MNIST test set.
@@ 6.2 Denoising performance
Here, we conducted denoising of the MNIST images across three distinct scenarios: noiseless, noisy, and real. To effectively demonstrate the performance of the QCAE, we introduced variations in Gaussian noise factors applied to the images because the MNIST does not carry noise. For each scenario, we tested 12 different configurations of Gaussian noise $\sigma=\{0.25,0.5,0.75,1\}$ across three environments (noiseless, noisy, real), and selected the combination that yielded the highest SSIM score for further experiments. The QCAE performance is evaluated for the denoising capability in both quantum simulations and actual machines.
Figure 10: Comparative analysis of the CCAE and specifically QCAE performance under varying conditions, including noiseless, noisy, and real scenarios, with noise factor variations represented by $\sigma=\{0.25,0.5,0.75,1\}$. Evaluation was conducted on the MNIST test set (digits 0 and 1 only).
To analyze the extent of QCAE denoising capabilities, we utilize it to restructure, denoise, and generate images from a set of Gaussian noisy inputs with varying noise factors values. Even in the presence of high levels of noise, characterized by a noise factor of $\sigma=1$, images that appear indistinguishable to the human eye are effectively processed by QCAE. The algorithm consistently delivers accurate and visually appealing results, demonstrating its robust performance in handling noisy images. Despite testing in noisy and real quantum machine environments, the image maintains readability and closely resembles the original, although with a slight degradation compared to a noiseless environment. As depicted in Fig. 11, denoising results with variations in Gaussian noise levels and quantum machine environments are shown for digit ‘0’ images selected from the same MNIST test subset (digits 0 and 1) used in prior evaluations. This ensures consistency across all noise levels and quantum environments.
@@ 6.3 Comparison of PQCs and PSR optimization
Figure 12: The SSIM value of $p$ ranges from 1 to 5 across all circuits, indicating that better solutions are associated with longer circuit layers. Evaluation was conducted on the same MNIST test subset (digits 0 and 1) used in earlier sections.
We also conducted comprehensive experiments on all circuits, both with and without PSR, resulting in a noteworthy reduction in SSIM values when tested without PSR, as shown in Table 1. Remarkably, our proposed circuit demonstrated the ability to draw comparisons with other circuits, even in the absence of PSR. One plausible explanation is that $R_{Z}$ gates can be implemented with greater efficiency compared to other single-qubit rotation operations [47]. As we observed, the circuits exhibit lower and significantly higher SSIM values with and without PSR, respectively. This underscores the significance of incorporating PSR optimization to enhance performance. It is noteworthy that the $p$ value significantly influences the results, underscoring its high impact on the outcomes. However, as the parameter $p$ increases, there is a corresponding growth in the circuit execution time. Additionally, our proposed method requires a minimum of two qubits for the efficient denoising of images. It also illustrates the progressive increase in SSIM values corresponding to the varying number of $p$ layers. As additional layers are added, the QAOA showcases the potential for enhanced approximations towards the optimal solution. Furthermore, it presents the results of experiments conducted on QAOA, both with and without PSR optimization. The findings emphasize the significance of employing this optimization technique to identify a lower-cost function, leading to higher accuracy. Through the utilization of the PSR optimization, there is a notable enhancement in SSIM value.